### v1

# Using quantum computers to identify prime numbers via entanglement dynamics

Abstract
The identification of prime numbers stands as a fundamental mathematical pursuit with significant historical and contemporary significance. Exploring potential connections between prime number theory and quantum physics represents a compelling frontier in scientific inquiry. In a recent study by A. L. M. Southier et al. [Phys. Rev. A 108, 042404 (2023)], the entanglement dynamics of two harmonic oscillators (or of two spin $s$ particles) initially prepared in a separable-coherent state was demonstrated to offer a pathway for prime number identification. Building upon this foundation, this Letter presents a generalized approach and outlines a deterministic algorithm for realizing this theoretical concept on qubit-based quantum computers. Our analysis reveals that the diagonal unitary operations employed in our algorithm exhibit a polynomial time complexity of degree two, contrasting with the previously reported exponential complexity of general diagonal unitaries [J. Welch et al., New J. Phys. 16, 033040 (2014)]. This advancement underscores the potential of quantum computing for prime number identification and related computational challenges.
Prime numbers; Quantum entanglement; Entanglement dynamics; Quantum Computation
Introduction.— The quest to reliably and efficiently identify prime numbers remains a topic of great interest in number theory Bressoud1989 ; crandall ; Baillie2021 ; Granville2005 , particularly due to its intriguing connection with the non-trivial zeros of Riemann’s zeta function schmayer ; wolf ; Feiler2013 ; Sierra2008 . Over the centuries, numerous classical algorithms have been devised for identifying primes, each offering its own set of advantages and limitations Bressoud1989 ; crandall ; aaronson . Among these, the AKS primality test stands out as the first deterministic algorithm to exhibit polynomial time complexity for verifying the primality of individual integers, albeit with a polynomial degree that renders it less efficient for larger numbers AKS . Conversely, the Sieve of Eratosthenes, dating back to ancient times, offers a simpler approach, focusing on identifying all prime numbers within a specified range $N$. Its time complexity, $\mathcal{O}(N\log\log N)$, renders it particularly efficient for this purpose, making it a stalwart method in the realm of prime number identification crandall .
While classical algorithms for prime number identification Sieve2017 ; Miller1976 ; Adleman1983 ; AKS have undergone significant development, their adaptation to the realm of quantum computing remains relatively limited, with only a few studies exploring this and related problems Donis-Vela2017 ; Chau1997 ; Li2012 . However, the intersection of such questions with experimental physics presents a promising avenue for the development of more intuitive quantum algorithms. A notable recent study alexandre proposed an innovative approach to primality testing using quantum optics. In their work, researchers devised an experiment involving the entanglement of two quantum harmonic oscillators initially prepared in coherent states, followed by the measurement of the reduced linear entropy of one of them. Intriguingly, they theorized that information regarding prime numbers could be extracted from the Fourier modes of the reduced linear entropy: prime numbers were expected to adhere to a lower bound curve, while composite numbers would consistently surpass this bound. Although the experimental implementation has yet to be realized, their theoretical groundwork has laid the foundation for us to generalize their approach and to develop a deterministic algorithm tailored for implementation on qubit-based quantum computers.
In this letter, we build upon the theoretical framework proposed in Ref. alexandre , aiming to adapt it for implementation on qubit-based quantum computers by removing certain restrictions imposed on the Hamiltonian and initial states. As a result, we demonstrate, as detailed in the Supplemental Material (SM), that the class of diagonal unitary gates utilized in our approach can be implemented in polynomial time, contrary to the expectations set forth in Refs. Bullock2004 ; welch . Our algorithm is designed to determine all prime numbers within a given range $N$ through the manipulation of a bipartite system $AB$ and the measurement of the linear entropy of entanglement Bennett1996 ; Vidal1999 ; Basso2022 of subsystem $A$ over a period $T$.
Our methodology unfolds in several steps. Firstly, we modify the definitions to align with the peculiarities of qubit-based quantum computing. Secondly, we select a suitable initial state that can be efficiently prepared. Thirdly, we efficiently prepare an evolved state $|\psi(t)\rangle_{AB}$ using the techniques outlined in Ref. welch , which surprisingly results in polynomial gate cost, marking an unexpected exponential cost reduction in comparison to the general case. Subsequently, we measure the reduced purity of subsystem $A$, a task that can be executed efficiently ekert . Following this, we calculate the Fourier modes of the reduced purity function via numerical integration methods Press2007 .
Given a dataset encompassing all points within the range $N$, our algorithm enables the deterministic identification of Fourier modes corresponding to prime numbers, allowing for the distinction between primes and composites. We quantify the number of gates utilized at each step, with a specific focus on $Z$-rotations, Controlled-NOT and Hadamard gates. Additionally, we discuss simulations conducted using Qiskit Qiskit and explore potential enhancements to our algorithm for more efficient implementation on real qubit-based quantum hardware.
Preliminary definitions.— We begin by establishing key definitions for our bipartite system $AB$. Let $A$ and $B$ represent the respective subsystems, each characterized by a time-independent Hamiltonian $\hat{H}_{A}$ and $\hat{H}_{B}$, where $\hat{H}_{A}=\hat{H}_{B}$. We define a bipartite Hamiltonian $\hat{H}_{AB}=\lambda\hat{H}_{A}\otimes\hat{H}_{B}$, with $\lambda\in\mathbb{R}$ denoting the coupling constant. The corresponding time-evolution operator is given by $\hat{U}(t)=e^{-i\hat{H}_{AB}t/\hbar}$ Nielsen2000 .
To ensure the distinction between prime and composite numbers, we employ an initial state $|\psi(0)\rangle_{AB}$ that is a product state of identical subsystem states, i.e., $|\psi(0)\rangle_{AB}=|\phi\rangle_{A}\otimes|\phi\rangle_{B}$. A suitable choice for these individual states is obtained by expressing them in the eigenbasis $\{|E_{n_{S}}\rangle\}_{n_{S}=1}^{d}$ of each respective subsystem’s Hamiltonian:
$|\phi\rangle_{S}=\sum_{n_{S}=1}^{d}c_{n_{S}}|E_{n_{S}}\rangle,$ (1)
where $S=A,B$ represents the subsystem index, $d$ is the dimension of each subsystem, and $c_{n_{S}}\neq 0$ are the initial state coefficients. If $\{E_{n_{S}}\}_{n_{S}=1}^{d}$ denote the eigenvalues of each subsystem Hamiltonian, the evolved state at time $t$ is obtained as $|\psi(t)\rangle_{AB}=\hat{U}(t)|\psi(0)\rangle_{AB}=\sum_{n_{A},n_{B}=1}^{d}c_{n_{A}}c_{n_{B}}e^{-i\lambda E_{n_{A}}E_{n_{B}}t/\hbar}|E_{n_{A}}E_{n_{B}}\rangle$.
Our main condition requires that the energy levels of both individual Hamiltonians are equidistant. This condition is expressed as $E_{n_{S}}=n_{S}\mu$ for some constant $\mu\in\mathbb{R}$. Defining $\omega=\lambda\mu^{2}/\hbar$, we find that
$\hat{U}(t)=\sum_{n_{A},n_{B}=1}^{d}e^{-i\omega n_{A}n_{B}t}|E_{n_{A}}E_{n_{B}}\rangle\langle E_{n_{A}}E_{n_{B}}|,$ (2)
and
$|\psi(t)\rangle_{AB}=\sum_{n_{A},n_{B}=1}^{d}c_{n_{A}}c_{n_{B}}e^{-i\omega n_{A}n_{B}t}|E_{n_{A}}E_{n_{B}}\rangle.$ (3)
A key result of our research is the demonstration of high gate-efficiency for implementing the diagonal unitary gate specified in Eq. (2), as detailed in Sec. I.3 of the SM. We show that the gate cost for constructing the $q$-qubit unitary gate $\hat{U}(t)$ using this method is a polynomial function $G_{2}(q)=\frac{3}{4}q^{2}+q$. This result not only facilitates prime number identification but also paves the way for efficient implementation of similar unitary gates in future qubit-based quantum computing research.
Reduced purity.— We designate subsystem $A$ for computing the reduced purity $\gamma_{A}(t)$ without loss of generality. Let us begin by revisiting the definition of the reduced density operator $\hat{\rho}_{A}(t)$ for a system $AB$ with density operator $\hat{\rho}_{AB}(t)$, given as $\hat{\rho}_{A}(t)=\operatorname{Tr}_{B}(\hat{\rho}_{AB}(t))$, where $\operatorname{Tr}_{B}(.)$ denotes the partial trace function Maziero2017 over subsystem $B$. The reduced purity function, $\gamma_{A}(t)=\operatorname{Tr}(\hat{\rho}^{2}_{A}(t))$, can then be computed from $\hat{\rho}_{A}(t)$. This quantity is related to the linear entanglement entropy by $E_{l}(|\psi(t)\rangle_{AB})=1-\gamma_{A}(t).$
Given our definition in Eq. (3), which implies that $\hat{\rho}_{AB}(t)$ represents a pure state, obtaining $\hat{\rho}_{A}(t)$ is straightforward. Subsequently, the reduced purity can be expressed as
$\gamma_{A}(t)=\sum_{j,k,l,m=1}^{d}|c_{j}|^{2}|c_{k}|^{2}|c_{l}|^{2}|c_{m}|^{2}e^{-i\omega t(j-k)(l-m)}.$ (4)
It is noteworthy to highlight several properties of the function $\gamma_{A}(t)$ defined in Eq. (4). Firstly, it exhibits time periodicity with period $T=2\pi/\omega$, a characteristic stemming directly from the time evolution of our system. Secondly, a notable observation arises from the structure of the sum in Eq. (4): the indices $j,k,l,$ and $m$ all take the same values. As a consequence, the imaginary parts of the phases $e^{-i\omega t(j-k)(l-m)}$ for a fixed $t$ mutually cancel each other. This cancellation is crucial, ensuring that $\gamma_{A}(t)$ remains a real-valued function.
Exploiting this property, we can demonstrate that $\gamma_{A}(t)$ is symmetric about half the period $T$. To illustrate, let $h$ denote some increment relative to $T/2=\pi/\omega$, and let $s=(j-k)(l-m)$. We find:
$e^{-i\omega(T/2+h)s}$ $=e^{-i\omega(\pi/\omega+h)s}$
$=\cos(\pi s+\omega hs)-i\sin(\pi s+\omega hn)$
$=\cos(\pi s)\cos(\omega hs)-i\sin(\pi s+\omega hn).$ (5)
Since the reduced purity function (4) only contains a real part, we can disregard the imaginary part in Eq. (Using quantum computers to identify prime numbers via entanglement dynamics). Furthermore, leveraging $\cos(\omega hs)=\cos(-\omega hs)$, we establish $\gamma_{A}(T/2+h)=\gamma_{A}(T/2-h)$, confirming the desired symmetry property. This symmetry property holds significance in our algorithm, as it enables us to halve the number of times we need to execute the quantum circuit to obtain $\gamma_{A}(t)$.
Mapping prime numbers with Fourier modes.— The reduced purity function given by Eq. (4) can be expressed as a finite sum of cosines, where the maximum number of Fourier modes $\alpha_{n}$ is $(d-1)^{2}$. Therefore, employing a Fourier expansion in this scenario yields:
$\gamma_{A}(t)=\alpha_{0}+\sum_{n=1}^{(d-1)^{2}}\alpha_{n}\cos(n\omega t),$ (6)
where $\alpha_{0}$ represents the average value and $\alpha_{n}$ are the Fourier modes Arfken2013 .
To compute the Fourier modes $\alpha_{n}$, we utilize the expression:
$\alpha_{n}=4\sum_{k,m=1}^{d-1}\sum_{j>k}^{d}\sum_{l>m}^{d}|c_{j}|^{2}|c_{k}|^{2}|c_{l}|^{2}|c_{m}|^{2}\delta^{n}_{(j-k)(l-m)},$ (7)
where $\delta^{n}_{(j-k)(l-m)}$ represents the Kronecker delta function ensuring the resonance condition for the Fourier modes. This formulation allows us to decompose the reduced purity $\gamma_{A}(t)$ into its constituent Fourier components, facilitating the identification of prime numbers based on their distinct Fourier signatures.
For prime $n$, the trivial decomposition is $(j-k)=n$ and $(l-m)=1$, and vice versa. This results in a unique decomposition that corresponds to the expected behavior for prime numbers. However, if $n$ is composite, it possesses non-trivial decompositions as well. To examine the impact of these decompositions on the Fourier modes expressed in Eq. (7), let us define the lower bound $B_{n}$ as the value obtained from Eq. (7) using the trivial decomposition of $n$. Hence, we have:
$B_{n}=8\sum_{k=1}^{d-n}\sum_{m=1}^{d-1}|c_{k}|^{2}|c_{m}|^{2}|c_{k+n}|^{2}|c_{m+1}|^{2}.$ (8)
This lower bound $B_{n}$ provides insight into the minimum value that the Fourier coefficient $\alpha_{n}$ can attain for a given composite $n$. Understanding this bound is crucial for discerning the distinct Fourier signatures associated with prime and composite numbers.
For $11\leq n\leq d-$ 1, we have $B_{n}>0$ as per Eq. (8). However, when $d\leq n\leq(d-1)^{2}$, the domain of $B_{n}$ can be extended such that $B_{n}=0$.
Now, let $\{y^{(n)}_{r}\}_{r=1}^{z}$ represent the sequence of $z$ distinct divisors of $n$ in increasing order of magnitude. Excluding the trivial cases $y^{(n)}_{1}=1$ and $y^{(n)}_{z}=n$, we find that in general:
$\alpha_{n}=B_{n}+4\sum_{r=2}^{z-1}\sum_{k=1}^{d-\frac{n}{y^{(n)}_{r}}}\sum_{m=1}^{d-y^{(n)}_{r}}|c_{k}|^{2}|c_{m}|^{2}|c_{k+\frac{n}{y^{(n)}_{r}}}|^{2}|c_{m+y^{(n)}_{r}}|^{2}.$ (9)
This expression for $\alpha_{n}$ encompasses both the contribution from the trivial decomposition and the contributions from the non-trivial divisors of $n$, enabling a comprehensive assessment of the Fourier modes associated with composite numbers.
In the domain $1\leq n\leq 2(d-1)$, we can confidently assert that $\alpha_{n}>B_{n}$ holds true. However, beyond this range, specifically in the interval $2(d-1)<n\leq(d-1)^{2}$, certain composite numbers $n=n_{0}$ may exhibit $\alpha_{n_{0}}=0$. This phenomenon arises because the first semiprimes (numbers that are the product of two prime numbers) are multiples of $2$. Consequently, for $n_{0}=2v$, where $v>d-1$ is a prime, there exist no values for the indices $k$ and $m$ in Eq. (9) that fall within their defined ranges in the summation.
As a result, we can confidently conclude that prime numbers always satisfy $\alpha_{n}=B_{n}$ in the interval $1\leq n\leq 2(d-1)$, where $B_{n}$ represents the extended lower bound defined in Eq. (8). Conversely, composite numbers consistently exhibit $\alpha_{n}>B_{n}$ within the same interval.
This insight provides a clear distinction between the Fourier modes associated with prime and composite numbers, aiding in the identification of prime numbers through their distinct Fourier signatures.
Here’s the summary of the expected values of $\alpha_{n}$ in the three regimes:
- Regime I: $11\leq n\leq d-$ 1. For prime numbers in this range, it holds true that $\alpha_{n}=B_{n}>0$; otherwise, $\alpha_{n}>B_{n}$.
- Regime II: $d\leq n\leq 2(d-1)$. Prime numbers in this interval exhibit $\alpha_{n}=B_{n}=0$, while composite numbers consistently demonstrate $\alpha_{n}>0$.
- Regime III: $2(d-1)<n\leq(d-1)^{2}$. Prime numbers within this regime always yield $\alpha_{n}=0$. However, some composite numbers may also yield $\alpha_{n}=0$ in this interval. Consequently, this regime cannot provide conclusive evidence regarding the primality of $n$.
This summary provides a clear delineation of the behavior of $\alpha_{n}$ across different regimes, aiding in the identification of prime numbers based on their Fourier modes.
Our regime of interest is $\mathcal{D}=\textbf{I}\cup\textbf{II}$. In $\mathcal{D}$, it is consistently true that:
$\alpha_{n}\geq B_{n},$ (10)
with equality achieved if and only if $n$ is a prime number. This inequality forms the cornerstone of our algorithm and serves as the basis for objectively distinguishing prime numbers from composites.
While our protocol enables the computation of $\alpha_{n}$, without knowledge of $B_{n}$ in Regime I, it is impossible to discern whether $\alpha_{n}=B_{n}$ or $\alpha_{n}\neq B_{n}$. A straightforward solution involves obtaining the analytical value of the lower bound $B_{n}$ within that regime, achievable by selecting a simple initial state and utilizing Eq. (8) subsequently. In our algorithm, for simplicity, we opt for an initial state of maximum superposition.
Quantum algorithm.— Below, we provide a structured description of all the steps necessary to develop our protocol. We also present here the number of gates necessary for each step.
1. Qubit Codification: To adapt our protocol to a qubit-based quantum computing algorithm, we need to adjust some of our definitions regarding the translation of qudits to qubits. Given that the bipartite system $AB$ has $d^{2}$ energy levels and we aim to utilize $q$ qubits instead of two qudits, the condition is imposed that:
$d^{2}=2^{q}.$ (11)
Equation (11) inherently assumes that $d$ is a power of $2$. However, in scenarios where $d$ is not a power of $2$, $q$ may not be an integer, rendering it irrelevant as the number of qubits. To accommodate such cases, we adjust our approach to determine the smallest power of $2$ greater than $d$. This can be mathematically expressed as finding $q$ such that $q=2\lceil\log_{2}(d)\rceil$, where $\lceil$.$\rceil$ . ⌉ denotes the ceiling function. Additionally, it follows from Eq. (11) that $q/2$ qubits will represent subsystem $A$ and $q/2$ qubits will represent subsystem $B$. Therefore, we conveniently assign the first half of qubits to represent $A$ and the remaining half to represent $B$.
2. Initial State Flexibility: The initial state $|\psi(0)\rangle_{AB}$ is defined as the product state $|\psi(0)\rangle_{AB}=|\phi\rangle_{A}\otimes|\phi\rangle_{B}$, where the coefficients $c_{n_{S}}$ of the subsystem states $|\phi\rangle_{S}=\sum_{n_{S}=1}^{d}c_{n_{S}}|E_{n_{S}}\rangle$ must satisfy $c_{n_{S}}\neq 0$. Leveraging this degree of freedom, we opt for convenience by employing an initial state that achieves maximum superposition, expressed as $|\psi(0)\rangle_{AB}=\frac{1}{d}\sum_{n_{A},n_{B}=1}^{d}|E_{n_{A}}E_{n_{B}}\rangle$. Here, we implicitly define the eigenbasis $\{E_{n_{S}}\}_{n_{S}=1}^{d}$ as the computational basis for each set of $q/2$ qubits. To produce this initial state, we apply a series of Hadamard gates $\hat{H}$ Nielsen2000 to all $q$ qubits:
$|\psi(0)\rangle_{AB}=\hat{H}^{\otimes q}|000...0\rangle.$ (12)
It is evident that the number of gates required here to generate this initial state is simply:
$G_{1}(q)=q.$ (13)
3. Evolved State Preparation: The detailed results regarding this item are provided in the SM. To obtain the evolved state $|\psi(t)\rangle_{AB}$ of Eq. (3), we employ the method outlined in Ref. welch to construct $\hat{U}(t)$, which is then applied to the initial state $|\psi(0)\rangle_{AB}$. While this method is generally expected to necessitate an exponential number of gates, as demonstrated in Sec. I.3, in our case implementing $\hat{U}(t)$ only requires a polynomial number of gates. Initially, we determine the $12^{q}-$ 1 Walsh angles $a_{j}(t)$ Walsh1923 ; Fine1949 ; Zhihua1983 ; Yuen1975 . By definition, the Walsh angles are expressed as
$a_{j}(t)=\frac{1}{2^{q}}\sum_{k=0}^{2^{q}-1}f_{k}(t)w_{jk},$ (14)
where $w_{jk}$ denotes the Paley-ordered discrete Walsh functions. Each coefficient $f_{k}(t)$ in Eq. (14) is derived from the unitary operator $\hat{U}(t)$. This operator is defined by its action on the eigenbasis $\{|k\rangle\}_{k=0}^{2^{q}-1}$ of $\hat{H}_{AB}$:
$\hat{U}(t)=e^{i\hat{f}(t)},\quad\text{where}\quad\hat{f}(t)=\sum_{k=0}^{2^{q}-1}f_{k}(t)|k\rangle\langle k|.$ (15)
Together with the Walsh angles $a_{j}(t)$, we can obtain $\hat{U}(t)$ using the formalism of Walsh operators $\hat{w}_{j}$. These operators are expressed as a tensor product of powers of Pauli $\hat{Z}$ operators. Specifically, we have $\hat{w}_{j}=\bigotimes_{i=1}^{q}(\hat{Z}_{i})^{j_{i}}$, where we define $(\hat{Z}_{i})^{1}=\hat{Z_{i}}$ and $(\hat{Z}_{i})^{0}=\hat{I}$, and $(j_{q}...j_{2}j_{1})$ represents the binary representation of the integer $j$, with the most significant non-zero bit (MSB) on the left. This yields
$\hat{U}(t)=\prod_{j=1}^{2^{q}-1}e^{ia_{j}(t)\hat{w}_{j}}.$ (16)
To produce the exponential operators $e^{ia_{j}(t)\hat{w_{j}}}$, we use the identity presented in Sec. I.1 of the SM, which allows us to express this exponential operator as a single $Z$-rotation $\hat{R}_{z}(\theta_{j}(t))$ on qubit $q_{m_{h_{j}}}$ and staircases of controlled-NOT gates targeted on $q_{m_{h_{j}}}$, where $1\leq m_{h_{j}}$ denotes the position of the MSB of $j$. These $\hat{R}_{z}(\theta_{j}(t))$ rotations have angles $\theta_{j}(t)$ related to $a_{j}(t)$ through $\theta_{j}(t)=-2a_{j}(t)$. According to the results shown in Sec. I.3 of the SM, most Walsh angles $a_{j}(t)$ for our $\hat{U}(t)$ are null. Moreover, the nonzero Walsh angles $a_{j}(t)$ follow simple formulas, eliminating the need for the matrix products of Eq. (14). Consequently, implementing $\hat{U}(t)$ requires only $\frac{3}{4}q^{2}+q$ gates. Therefore, preparing $|\psi(t)\rangle_{AB}$ from the initial state $|\psi(0)\rangle_{AB}$ demands a number of gates given by
$G_{2}(q)=\frac{3}{4}q^{2}+q.$ (17)
4. Reduced Purity Estimation. This step involves efficiently obtaining the reduced purity $\gamma_{A}(t)$ of Eq. (4) by utilizing techniques from Ref. ekert . The quantum circuit employed here resembles the SWAP test circuit buhrman ; barenco and employs an ancilla qubit $q_{0}$ and two copies of $q$ qubits prepared in the same pure state. The operations sequence for this quantum circuit is as follows: a Hadamard gate on $q_{0}$, qubit-qubit controlled-SWAP gates between the first $q/2$ qubits of each copy, with $q_{0}$ as the control qubit, another Hadamard gate on $q_{0}$ and a measurement of $q_{0}$ in the computational basis. After repeatedly executing the circuit, we estimate the probability $P_{0}$ of obtaining the state $|0\rangle$ for $q_{0}$. Then, as detailed in Sec. II of the SM, the reduced purity over time can be estimated using the expression $\gamma_{A}(t)=2P_{0}(t)-1$.
Since this quantum circuit entails two Hadamard gates on the ancilla qubit $q_{0}$ and $q/2$ controlled-SWAP gates between the other $2q$ qubits, where each controlled-SWAP gate consists of three controlled-NOT gates, this step involves a total number of gates given by
$G_{3}(q)=\frac{3}{2}q+2.$ (18)
5. Fourier Modes Calculation. With all the previous ingredients in hand, we can compute the Fourier modes of Eq. (7) via numerical integration of the reduced purity $\gamma_{A}(t)$, which is obtained as outlined in the previous step. In Regime I, we obtain the lower bound $B_{n}$ of Eq. (8) using the initial state of Eq. (12). In this case, $c_{j}=\frac{1}{d}$ for any $j$, and the corresponding lower bound $B_{n}$ interpolation in this range of $n$ is a straight line with a negative slope. In Regime II, the lower bound is $B_{n}=0$. The expression for $B_{n}$ in the regime of interest $\mathcal{D}=$ I $\cup$ II can then be written as
$B_{n}=\begin{cases}\frac{-8(d-1)}{d^{4}}n+\frac{8d-8}{d^{3}}&\text{if }n\in\textbf{I},\\ 0&\text{if }n\in\textbf{II}.\end{cases}$
Considering the remarks made in the previous section, we know that in a graph of Fourier modes, every prime number must have a corresponding $\alpha_{n}$ position belonging exactly to the interpolated curve of $B_{n}$. Any composite number in the regime of interest $\mathcal{D}$ has $\alpha_{n}>B_{n}$ and thus is necessarily above $B_{n}$.
Now, using Fourier analysis, the Fourier modes $\alpha_{n}$ are calculated by the integral
[…]
Normally, Eq. (19) would be an integral over the whole period $T$, but we are employing the property of the symmetry of $\gamma_{A}(t)$, presented earlier in this letter. After calculating the Fourier modes $\alpha_{n}$, the last part of our algorithm involves comparing the value of $\alpha_{n}$ with the analytical lower bound $B_{n}$. In this final step, the numerical integration is done in $p$ partitions, resulting in an equivalent number of points used for $\gamma_{A}(t)$ in the interval $20\leq t\leq T/$ 2. Consequently, to achieve a desired precision $\epsilon$, our quantum circuit requires at least $p$ executions. Currently, the exact scaling of $p$ with respect to $d$, for a given $\epsilon$, remains undetermined.
Figure 1: Schematic of the quantum circuit for our algorithm for $d=4$. Each colored box corresponds to a stage of the circuit and represents the set of operations in the respective step for the two copies of $q$ qubits. In the first stage, we prepare the initial state of maximum superposition $|\psi(0)\rangle_{AB}$ for both copies of $q=4$ qubits, starting from the state $|0\rangle$ for each qubit. The second stage is used for the efficient state preparation of $|\psi(t)\rangle_{AB}$ for each copy, where the total number of gates, including both copies, is given by $\frac{3}{2}q^{2}+2q$. In the last stage, with the aid of an ancilla qubit $q_{0}$, we apply the gates corresponding to the variation of the SWAP test to extract the reduced purity $\gamma_{A}(t)$ after executing the circuit several times.
Simulations.— To evaluate the applicability of our algorithm, simulations were performed using IBM’s Qiskit framework (version 0.45.1). These simulations targeted three distinct values of $d$, with results depicted in blue in Fig. 2. For all the simulations, we used $10^{5}$ shots and fixed $\omega=0.1\,\text{s}^{-1}$, as changing the value of $\omega$ has no effect on the Fourier modes $\alpha_{n}$. Regarding the number of executions $p$ of the circuit, we selected $p=375$, $p=1500$, and $p=6000$ for the dimensions $d=16$, $d=32$, and $d=64$, respectively. The values of $p$ were chosen to achieve roughly the same accuracy for the three values of $d$. Using Python (version 3.11.3) with the Scipy library (version 1.11.3), Fourier modes $\alpha_{n}$ were calculated with Simpson’s rule for the numerical integration of Eq. (19). Due to the substantial size of the quantum circuit for the three dimensions analyzed in our simulations, we present the circuit for a lower dimension, $d=4$, purely for illustrative purposes. This simplified example is shown in Fig. 1, allowing us to convey the structure without the complexity of the larger dimensions.
Figure 2: Comparison of simulation results with theoretical predictions for the Fourier modes of the reduced purity across different dimensions $(d=16,32,64)$. Red points represent the analytical values of $\alpha_{n}$, calculated directly from their theoretical expressions, while the green line stands for the minimum value for the Fourier modes, also derived from theoretical calculations. Blue points illustrate the Fourier modes obtained through numerical integration of the reduced purity $\gamma_{A}(t)$ extracted from the classical emulation of our quantum circuit. The numerical integration was performed using various partition values $(p=375,1500,6000)$. Prime numbers are expected to align with the lower bound $B_{n}$, whereas composite numbers appear above. Points closely approaching the lower bound $B_{n}$, yet not precisely on it, may be considered prime candidates, subject to verification due to the presence of statistical error.
Conclusions.— This work presents a qubit-based quantum algorithm for prime number identification, rooted in the analysis of entangled subsystem dynamics. By strategically selecting a bipartite Hamiltonian of dimension $d^{2}$ characterized by energy quantization, we demonstrate that the Fourier modes $\alpha_{n}$ of the reduced purity display distinctive patterns. These patterns enable the differentiation between prime and composite numbers for any integer $n$ such that $1\leq n\leq 2(d-1)$. To implement this protocol as an algorithm for qubit-based quantum computing, we map this $2$-qudit system into a $q$-qubit system using the expression $d^{2}=2^{q}$. In this mapping, as an independent result, we found that the local unitary gates with the form of Eq. (2) may be implemented in polynomial time using the method of Ref. welch .
Executing our quantum circuit involves three stages, as illustrated in Fig. 1: preparation of the initial state $|\psi(0)\rangle_{AB}$ for both sets of $q$ qubits, preparation of the evolved state $|\psi(t)\rangle_{AB}$ also for both sets, and estimating the reduced purity $\gamma_{A}(t)$. These stages have respective gate costs: $2G_{1}=2q$, $2G_{2}(q)=\frac{3}{2}q^{2}+q$, and $G_{3}=\frac{3}{2}q+2$. In particular, $G_{2}(q)$ represents the number of gates necessary to implement local unitary gates that share the same mathematical structure as $\hat{U}(t)$, where this implementation is accomplished using the methodology outlined in Ref. welch . Contrary to the anticipated exponential cost of the method delineated in that work, we have demonstrated that $G_{2}(q)$ scales polynomially with the number of qubits $q$, thereby affirming the high efficiency of that method in preparing the class of states $|\psi(t)\rangle_{AB}$, as defined by Eq. (3). Thus, the total gate cost $G(q)$ for one execution of our quantum circuit is given by $G(q)=\frac{3}{2}q^{2}+\frac{11}{2}q+2$. By defining $N=2(d-1)$ and using $2^{q}=d^{2}$, we find that the asymptotic gate cost is $\mathcal{O}(\log^{2}_{10}(N))$, indicating a quadratic scaling in the number of digits of $N$.
Naturally, although this work presented ideal classical simulations, our algorithm could be implemented on quantum hardware. One major challenge that may arise, however, is that controlled-NOT and SWAP gates require good connectivity between the qubits, which is not always available in existing superconducting qubit-based quantum computers. A simple solution is to explore alternative quantum computing technologies. Trapped ion quantum computers, exemplified by those developed by IonQ and Quantinuum, offer extended coherence times and all-to-all connectivity between qubits, allowing for more effective implementation of these gates. Another approach would be to individually modify the last two stages. For the second stage, which involves preparing $|\psi(t)\rangle_{AB}$ for both copies, alternative methods could be considered, although we suspect that Ref. welch provides the most efficient technique for this purpose. For the third and final stage, since the reduced purity $\gamma_{A}(t)$ is related to the reduced density operator $\hat{\rho}_{A}(t)$, we could replace the direct measurement of $\gamma_{A}(t)$ with a state tomography of $\hat{\rho}_{A}(t)$. Modern tomographic methods, such as classical shadows tomography Aaronson2018 ; Huang2020 ; Nguyen2022 ; Bu2024 , show promise and can be employed for the estimation of $\gamma_{A}(t)$, requiring only $q$ qubits instead of $12q+$ 1.
Importantly, our demonstration that the unitary gate $\hat{U}(t)$ can be implemented with a polynomial number of gates, contrary to the general expectation of exponential gate requirements, paves the way for the efficient realization of similar transformations in quantum computing. This finding not only enhances the feasibility of executing our prime number identification algorithm on actual quantum systems but also suggests a broader applicability for optimizing other quantum algorithms. As quantum computing continues to evolve, the adaptation of such efficient unitary operations will be crucial in overcoming the limitations posed by gate complexities and in harnessing the full potential of quantum technologies. With advancements in fault tolerance and increased access to a larger number of qubits in the future, our algorithm holds the potential to verify the primality of significantly larger primes, marking a significant step forward in the practical application of quantum algorithms for fundamental computational problems.
Acknowledgements.
This work was supported by the Coordination for the Improvement of Higher Education Personnel (CAPES), Grant No. 23081.031640/2023-17, by the National Council for Scientific and Technological Development (CNPq), Grants No. 309862/2021-3, No. 409673/2022-6, and No. 421792/2022-1, and the National Institute for the Science and Technology of Quantum Information (INCT-IQ), Grant No. 465469/2014-0.